OpenStudy (anonymous):

You are instructed to create 500. mL of a 0.62 M phosphate buffer with a pH of 6.4. You have phosphoric acid and the sodium salts NaH2PO4, Na2HPO4, and Na3PO4 available. (Enter all numerical answers to three significant figures.) H3PO4(s) + H2O(l) equilibrium reaction arrow H3O+(aq) + H2PO4−(aq) Ka1 = 6.9 ✕ 10−3 H2PO4−(aq) + H2O(l) equilibrium reaction arrow H3O+(aq) + HPO42−(aq) Ka2 = 6.2 ✕ 10−8 HPO42−(aq) + H2O(l) equilibrium reaction arrow H3O+(aq) + PO43−(aq) Ka3 = 4.8 ✕ 10−13

OpenStudy (aaronq):

which 2 main species would you use for the buffer at pH 6.4?

OpenStudy (anonymous):

NaH2PO4 for acid component and Na2HPO4 for base component

OpenStudy (aaronq):

so now apply the henderson-hasselbalch equation \(pH=pKa+log\dfrac{[Na_2HPO_4]}{[NaH_2PO_4]}\) find the ratio of \(\dfrac{[Na_2HPO_4]}{[NaH_2PO_4]}\) then find the composition of each species in the buffer. you can use this as a reference http://www.xula.edu/chemistry/documents/biolab/Carroll%20lab%20Chap%203.pdf